Grim Company applies overhead on the basis of machine hours. Given the following data, compute overhead applied and the under- or overapplication of overhead for the period:
Estimated annual overhead cost   $3,400,000
Actual annual overhead cost   $3,370,000
Estimated machine hours   400,000
Actual machine hours   395,000
Question 5 options:
  
$3,357,500 applied and $12,500 overapplied.
  
$3,357,500 applied and $12,500 underapplied.
  
$3,370,000 applied and neither under- nor overapplied.
  
$3,400,000 applied and $12,500 overapplied.

Thank you
Answer #1

Solution:

Overhead application rate = Estimated overhead / estimated machine hours = 3400000/ 400000 = $8.5 per MH

Overhead applied = Actual machine hours * Overhead application rate = 395000* $8.5 = $3,357,500

Underapplied overhead = Actual overhead - applied overhead = 3370000 - 3357500 = $12,500

Hence "$3,357,500 applied and $12,500 underapplied" is correct.